Select the correct answer.

A software designer is mapping the streets for a new racing game. All of the streets are depicted as either perpendicular or parallel lines. The equation of the lane passing through A and B is -7x + 3y = -21.5. What is the equation of the central street PQ?

A. -3x + 4y = 3

B. -1.5x − 3.5y = -31.5

C. 2x + y = 20

D. -2.25x + y = -9.75


1
Expert's answer
2021-07-15T04:29:06-0400

The equation of the lane passing through "A" and "B," is


"-7x+3y=-21.5"

Solve for "y"


"y=\dfrac{7}{3}x-\dfrac{43}{6}"

Hence

"slope_1=\dfrac{7}{3}"

Since the central street "PQ" is perpendicular to the lane passing through "A" and "B," then


"slope_2=\dfrac{-1}{slope_1}=-\dfrac{3}{7}"

Point "P(7, 6)"


"6=-\dfrac{3}{7}(7)+b=>b=9"

The equation of the central street "PQ" is


"y=-\dfrac{3}{7}x+9"

"-1.5x-3.5y=-31.5"

B. -1.5x − 3.5y = -31.5